Write the equation of a line in slope-intercept form that is parallel to the line y= 7/8x-7 and containing point (8, 0).
Answer by lwsshak3(11628)

Write the equation of a line in slope-intercept form that is parallel to the line y= 7/8x-7 and containing point (8, 0)
***
Standard form for a straight line: y=mx+b, m=slope, b=y-intercept
y= 7x/8-7
slope of given line: m=7/8
slope of line parallel to given line has the same slope=7/8
Equation:
y=mx+b
solving for b using given point (8,0)
0=7*8/8+b
b=-7
Equation: y=7x/8-7
note: both lines are the same